Use the method of substitution to solve the system of linear equations.
2x-y=5
5x-2y=13
The solution is ?

Answer by tommyt3rd(5050)

1) 2x-y=5
2) 5x-2y=13


5x-2(2x-5)=13
x=3
y=2(3)-5=1